OpenStudy (anonymous):

Points U, V, and W are plotted on a number line. The coordinate of point V is –2. Points U and W are the same distance from point V. What could be the coordinates of U and W? A. U = –11, W = 7 B. U = –8, W = –6 C. U = –6, W = 6 D. U = 0, W = 2

OpenStudy (anonymous):

i will give metal

OpenStudy (anonymous):

Distance between U and V is |U - V| Distance between W and V is |W - V|

OpenStudy (anonymous):

A. V = -2, U = -11, W = 7 Distance between U and V is |U - V| = |-11 - (-2)| = 9 Distance between W and V is |W - V| = |7 - (-2)| = 9

OpenStudy (anonymous):

thx